# Online model learning with data-assimilated reservoir computers

### v1

@@ 2.1 Non-intrusive reduced-order model: POD-ESN
Figure 1: Streamwise velocity field of cylinder flow. Snapshots of (a) the truth; (b) the noisy data, which are the observable fields; (c) the POD reconstruction with $N_{m}=4$; and the instantaneous root mean squared (RMS) error field of panel (c) with respect to (d) the noisy data, and (e) the truth. The circles in (c) indicate the sensors, which are selected and numbered by the QR decomposition.
@@
$({\bm{\mathsf{R}}}{\bm{\mathsf{R}}}^{\mathrm{T}}+\lambda_{T}\mathbb{I}){\bm{\mathsf{W}}}_{\mathrm{out}}^{\mathrm{T}}={\bm{\mathsf{R}}}{\bm{\mathsf{U}}}^{\mathrm{T}}_{\mathrm{train}},$ (3)
where ${\bm{\mathsf{U}}}_{\mathrm{train}}$ is the training dataset, ${\bm{\mathsf{R}}}$ are the reservoir states corresponding to the training set, $\mathbb{I}$ is the identity matrix, and $\lambda_{T}$ is the Tikhonov regularization factor. The hyperparameters $\rho$ and $\sigma_{\mathrm{in}}$ are optimized with a recycle validation and Bayesian optimization. (For further details on the ESN training see [14, 9].)

### v2

@@ 2.1 Non-intrusive reduced-order model: POD-ESN
Figure 1: Streamwise velocity field of cylinder flow. Snapshots of (a) the truth; (b) the noisy data, which are the observable fields; (c) the POD reconstruction with $N_{m}=4$; and the instantaneous root mean squared (RMS) error field of panel (c) with respect to (d) the noisy data, and (e) the truth. The circles in (b) indicate the sensors, which are selected and numbered by the QR decomposition.
@@
$({\bm{\mathsf{R}}}{\bm{\mathsf{R}}}^{\mathrm{T}}+\lambda_{T}\mathbb{I}){\bm{\mathsf{W}}}_{\mathrm{out}}^{\mathrm{T}}={\bm{\mathsf{R}}}{\bm{\mathsf{\Phi}}}^{\mathrm{T}}_{\mathrm{train}},$ (3)
where ${\bm{\mathsf{\Phi}}}_{\mathrm{train}}$ is the training dataset, ${\bm{\mathsf{R}}}$ are the reservoir states corresponding to the training set, $\mathbb{I}$ is the identity matrix, and $\lambda_{T}$ is the Tikhonov regularization factor. The hyperparameters $\rho$ and $\sigma_{\mathrm{in}}$ are optimized with a recycle validation and Bayesian optimization. (For further details on the ESN training see [14, 9].)